x-3z = -2
3x + y - 2z = 5
2x + 2y + z = 4
could some one help me please

Answer by algebrapro18(206)

You can either solve the system by back substitution or elimination. I chose to solve it by substitution.
So we solve the first equation for x and get x = -2+3z. Now we can substitute that into the other equations. I'll leave the algebra for you to do but you will end up with:
y+7z=11
2y+7z=8
And now we can solve the system of 2 equations and 2 unknowns. Again I chose to do this by substitution but elimination would work as well. I solved the top equation for y and got y=11-7z. Then plugged that into the bottom equation so I could solve for z. Again I'll leave the algebra for you to do but you should get z = 2. Then you substitute that into your equation for y and end up with y = -3. Then the value in for z into the equation for x and you will see that x = 4.
So your final solution is (4,-3,2).